OpenStudy (anonymous):

1. Solve. y – 5 6 (1 point) y ≤ 1 y ≤ 11 y ≤ –11 y ≤ 30 2. Solve. 8 > d – 2 (1 point) d < 10 d > 10 d < 6 d > 4 3. Solve. 3.8 ≥ b + 4 (1 point) b ≤ –0.2 b ≤ 0.2 b ≤ 7.8 b ≤ 1.8

OpenStudy (anonymous):

y-5=6?

OpenStudy (anonymous):

Those are all straight forward. \[y \le11\] \[10>d\] \[b \le-.2\]

OpenStudy (anonymous):

thnxx

OpenStudy (anonymous):

8 > d - 2 If we add 2 to both sides, we have d. 8 + 2 > d - 2 + 2 10 > d which can be interpreted as d < 10, because of division by negatives flipping the inequality. 3.8 >= b + 4 Subtract 4 from both sides. 3.8 - 4 >= b + 4 -4 -.2 >= b If you want it the other way, b <= -.2 Hope this helps.

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

No problem.